# Mind the Gap Why Neural Memory Fails Under Semantic Density

### v1

@@ 5.6 Experiment 6: Cost Analysis
Beyond accuracy, we analyze the economic viability of context-based versus selective retrieval, a consideration that determines whether a memory architecture can be deployed at enterprise scale.
@@
We computed monthly costs using Claude Sonnet 4 pricing (\$3 per million input tokens, \$15 per million output tokens), assuming 50 output tokens per response and query volumes representative of realistic deployments.
@@
5.7 Experiment 7: Production Memory Failure Modes
This experiment moves beyond synthetic benchmarks to evaluate how production memory systems actually fail in realistic use cases, testing frontier model memory capabilities against four potential failure modes that users might encounter.
@@ 5.7 Experiment 7: Production Memory Failure Modes
We designed four tests targeting distinct failure modes. All API calls used temperature $t=0.0$ and top_p =1.0=1.0 to maximize reproducibility; we note that schema drift persists even at zero temperature, indicating it arises from the model’s learned priors rather than sampling stochasticity. For Claude models, we used the Anthropic API with model identifiers claude-sonnet-4-5-20250929; for GPT models, we used the OpenAI API with identifiers gpt-5.2 and gpt-4o-mini. The extraction prompt was: “Extract facts from the statement as JSON. Use this exact format: {“entity”: “…”, “predicate”: “…”, “value”: “…”}. Respond with ONLY the JSON, no other text.” No function calling or structured output mode was used; we tested the models’ natural extraction behavior.
For schema consistency, we provided 10 facts in natural language and asked the model to extract them as structured subject-predicate-object triples, measuring whether the model used consistent predicate names across extractions (for example, always using birth_place rather than varying between birth_place, was_born_in, and place_of_birth). For correction handling, we provided a sequence of statements where facts were updated (for example, “Alice’s phone is 555-1111” followed later by “Alice’s phone changed to 555-2222”), then queried for the current value, measuring whether the model returned only the latest value or ambiguously mentioned both old and new values. For retrieval accuracy, we loaded 20 facts into context and queried each, measuring simple retrieval performance. For positional bias, we varied the position of a target fact within a 50-fact context to test whether retrieval accuracy depended on where facts appeared in the prompt.
@@
When facts were corrected over the course of a conversation, Claude Sonnet 4.5 returned only the latest value 0% of the time—in all cases, it mentioned both old and new values in its response (for example, “Alice’s phone was 555-1111, but it changed to 555-2222”). While this behavior might seem helpful by providing context, it creates ambiguity for downstream systems and users who need definitive current values. If a user asks “What is Alice’s phone number?” they typically want the current number, not a history lesson. Knowledge Objects address this through explicit version chains that ensure corrections supersede prior values: queries return only the current version by default, with history available only when explicitly requested.
@@
At moderate scale ($N\leq 100$), all models achieved perfect retrieval accuracy with no detectable positional bias—facts at the beginning, middle, and end of context were retrieved equally well. This confirms that the Stability Gap we observed in Experiment 1 affects neural memory (gradient-based storage in weights) rather than context-window attention at small $N$. The attention mechanism successfully retrieves facts from context when $N$ is small; the problems arise from schema drift and version ambiguity, not retrieval failure.
@@
The failures we observe are not specific to any single model but reflect architectural limitations shared across frontier models from both Anthropic and OpenAI. All tested models exhibit schema drift, with each model producing different variants: birth_place becomes was_born_in (Claude), place_of_birth (GPT-5.2), or born_in (GPT-4o-mini). Interestingly, the GPT models handle corrections cleanly (100% returning only the latest value) while Claude consistently produces ambiguous responses mentioning both old and new values (0% clean corrections). This divergence suggests different internal approaches to representing and reasoning about temporal information across model families, though neither approach achieves the schema consistency required for reliable structured storage. Knowledge Objects address both problems through controlled vocabularies that constrain predicate choices during extraction and explicit version chains that ensure corrections cleanly supersede prior values without requiring the model to “remember” to omit old information.
@@ 6 Discussion
Neural memory systems that store facts in shared continuous parameters will experience interference proportional to semantic density, and only systems employing discrete addresses can guarantee zero interference between stored facts. This finding does not represent a limitation of current models that might be resolved through scaling or architectural refinements—rather, it reflects a geometric constraint inherent to continuous vector storage that no amount of engineering can circumvent. When two embedding vectors occupy nearby regions of the space, any linear combination that retrieves one will partially retrieve the other; this is not a bug but a mathematical property of inner products.
@@
Current AI systems attempt to accomplish everything within the “neocortex”—that is, neural weights and attention mechanisms—which explains why they fail at precise factual storage despite succeeding at pattern recognition, generation, and reasoning. Our Knowledge Objects implement the missing “hippocampus” in this analogy: not as a biological simulation, but as the engineering solution to the same computational problem that evolution solved through anatomical specialization.
@@
Relation to Production Systems.
Production AI assistants have empirically converged on discrete external storage for cross-session memory, validating the necessity of our approach. When users enable “memory” features in commercial systems, facts are stored in external databases rather than encoded into model weights. However, current implementations typically store unstructured text chunks such as “User mentioned they like Python” rather than structured triples such as (user, preferred_language, Python). This distinction matters significantly: structured storage enables deterministic lookup by subject and predicate, surgical updates that change single facts without affecting others, and version chains that track fact evolution over time. Our multi-model comparison demonstrates that schema drift and version ambiguity affect both Claude and GPT model families, confirming that these are architectural limitations of the unstructured approach rather than bugs in any particular model’s implementation.
@@
Our analysis has several limitations that should inform interpretation of the results. First, our neural baseline (LAM) is a simplified proxy for production systems that store facts via superposition; we treat it as a mechanistic lower bound that isolates the interference mechanism arising whenever semantically overlapping keys are stored in shared continuous parameters. Production implementations may include mitigations such as gating and regularization that shift the collapse curve rightward, but absent explicit pattern separation, overlap-driven interference persists when keys are semantically clustered—our claim concerns the existence of interference under semantic density, not the performance of any specific implementation. Second, our experiments use a single embedding model (all-MiniLM-L6-v2 with $d=384$ dimensions); different embedding dimensions would shift the quantitative results, though the qualitative finding—that collapse occurs at $N$ far below theoretical capacity—should hold. Third, we measure retrieval accuracy on synthetic fact corpora, not end-to-end task performance on downstream applications, though retrieval accuracy is a prerequisite for task success. Fourth, our multi-model comparison tests only three models from two providers; additional models might reveal different failure patterns, though the consistency we observe suggests these are general phenomena.

### v2

@@ 5.6 Experiment 6: Cost Analysis
Beyond accuracy, we analyze the economic viability of context-based versus selective retrieval. Production memory systems already use selective retrieval; this analysis establishes why context-based approaches are unviable, confirming that the industry convergence on selective retrieval was economically necessary rather than merely convenient.
@@
We computed monthly costs using Claude Sonnet 4 pricing as of January 2026 (\$3 per million input tokens, \$15 per million output tokens), assuming 50 output tokens per response and query volumes representative of realistic deployments. While absolute prices change as providers adjust rates, the scaling ratio ($O(N)$ vs $O(1)$) is determined by architecture, not vendor pricing.
@@
The Metabolic Analogy.
This cost differential mirrors a biological constraint. The human brain operates on approximately 20 watts yet stores a lifetime of episodic memories. It achieves this through sparse activation: recalling a specific memory activates only the neural populations encoding that memory, not the entire cortex. Context-based memory is analogous to a brain that must fire every neuron (attend to every token) to recall a single phone number, metabolically and computationally untenable. Knowledge Objects implement the biological strategy: retrieving one fact activates only the computational path for that fact ($O(1)$), not the entire knowledge base ($O(N)$). The Orthogonality Constraint may thus reflect not just a geometric limitation but nature’s solution to the energy efficiency requirements of scalable episodic memory.
5.7 Experiment 7: Model Extraction Failure Modes
This experiment evaluates the underlying models used in production memory systems, testing their ability to maintain schema consistency and handle corrections. We test the models (Claude, GPT) via API rather than the production features (ChatGPT Memory, Claude Memory), which may include additional engineering such as structured output modes, fine-tuned extractors, or post-processing. Our goal is to isolate the architectural limitations of generative extraction that any LLM-based memory system must contend with.
@@ 5.7 Experiment 7: Model Extraction Failure Modes
We designed four tests targeting distinct failure modes. All API calls used temperature $t=0.0$ and top_p =1.0=1.0 to maximize reproducibility; we note that schema drift persists even at zero temperature, indicating it arises from the model’s learned priors rather than sampling stochasticity. For Claude models, we used the Anthropic API with model identifier claude-sonnet-4-5-20250929; for GPT models, we used the OpenAI API with identifiers gpt-5.2 and gpt-4o-mini (model identifiers as of January 2026). The extraction prompt was: “Extract facts from the statement as JSON. Use this exact format: {“entity”: “…”, “predicate”: “…”, “value”: “…”}. Respond with ONLY the JSON, no other text.” No function calling or structured output mode was used; we tested the models’ natural extraction behavior.
For schema consistency, we provided 10 facts in natural language and asked the model to extract them as structured subject-predicate-object facts, measuring whether the model used consistent predicate names across extractions (for example, always using birth_place rather than varying between birth_place, was_born_in, and place_of_birth). For correction handling, we provided a sequence of statements where facts were updated (for example, “Alice’s phone is 555-1111” followed later by “Alice’s phone changed to 555-2222”), then queried for the current value, measuring whether the model returned only the latest value or ambiguously mentioned both old and new values. For retrieval accuracy, we loaded 20 facts into context and queried each, measuring simple retrieval performance. For positional bias, we varied the position of a target fact within a 50-fact context to test whether retrieval accuracy depended on where facts appeared in the prompt.
@@
When facts were corrected over the course of a conversation, Claude Sonnet 4.5 returned only the latest value 0% of the time, in all cases, it mentioned both old and new values in its response (for example, “Alice’s phone was 555-1111, but it changed to 555-2222”). While this behavior might seem helpful by providing context, it creates ambiguity for downstream systems and users who need definitive current values. If a user asks “What is Alice’s phone number?” they typically want the current number, not a history lesson. Knowledge Objects address this through explicit version chains that ensure corrections supersede prior values: queries return only the current version by default, with history available only when explicitly requested.
@@
At moderate scale ($N\leq 100$), all models achieved perfect retrieval accuracy with no detectable positional bias, facts at the beginning, middle, and end of context were retrieved equally well. This confirms that the Semantic Interference we observed in Experiment 1 affects neural memory (gradient-based storage in weights) rather than context-window attention at small $N$. The attention mechanism successfully retrieves facts from context when $N$ is small; the problems arise from schema drift and version ambiguity, not retrieval failure.
@@
System-Level Reliability Analysis.
While 70% per-extraction schema consistency might appear acceptable for individual queries, we can model the system-level implications using a Markov chain analysis. If each extraction has independent probability $p_{\text{drift}}$ of introducing a schema error, then after $T$ extractions the probability of maintaining a completely consistent schema is $(1-p_{\text{drift}})^{T}$. For Claude at $p_{\text{drift}}=0.30$, system reliability drops to 50% after just $T\approx 2$ extractions and to 10% after $T\approx 7$ extractions. For GPT-4o-mini at $p_{\text{drift}}=0.45$, 10% reliability is reached after only $T\approx 4$ extractions. This analysis reveals that “passable” per-extraction accuracy translates to catastrophic system-level failure: a memory system handling dozens of facts per user would experience near-certain schema corruption. Knowledge Objects with controlled vocabularies achieve $p_{\text{drift}}<0.001$ by construction, maintaining 95% system reliability past $T=50$ extractions. Figure 6 visualizes this exponential decay, showing that even the best-performing model (Claude at 70%) drops below coin-flip reliability after just 2 extractions.
Figure 6: Schema Reliability Decays Exponentially. Probability of maintaining a completely consistent schema after $T$ fact extractions, modeled as $(1-p_{\text{drift}})^{T}$ where $p_{\text{drift}}$ is the per-extraction schema error rate. Claude (70% per-extraction accuracy) drops to 50% system reliability after just 2 extractions; GPT-4o-mini (55%) drops below 10% after 4 extractions. Knowledge Objects with controlled vocabularies maintain near-perfect reliability by enforcing schema consistency at write time. The green dashed line indicates a 95% enterprise SLA; only controlled vocabularies can meet this threshold beyond a handful of extractions.
@@
The failures we observe reflect architectural limitations of generative extraction shared across frontier models from both Anthropic and OpenAI. All tested models exhibit schema drift, with each model producing different variants: birth_place becomes was_born_in (Claude), place_of_birth (GPT-5.2), or born_in (GPT-4o-mini). Interestingly, the GPT models handle corrections cleanly (100% returning only the latest value) while Claude consistently produces ambiguous responses mentioning both old and new values (0% clean corrections). This divergence suggests different internal approaches to representing and reasoning about temporal information across model families.
Relationship to Production Features.
We tested the underlying models via API, not the production memory features (ChatGPT Memory, Claude Memory) which may include additional mitigations such as structured output modes, function calling with schemas, or fine-tuned extraction models. However, our finding that schema drift persists even within single sessions, where models can observe their own prior extractions, suggests the problem is fundamental to generative extraction under weak constraints, not merely a cross-session consistency issue that engineering could easily solve.
A natural objection is that function-calling or structured output modes with strict schemas would eliminate drift entirely. This is partially true: if you define birth_place as the only valid predicate for birthplace facts, the model cannot drift to was_born_in. However, this shifts rather than solves the problem. In open-ended agent scenarios, the schema is not known a priori, the system must decide which predicates to create as it encounters new information. When the model invents predicates on-the-fly (e.g., encountering “User prefers dark mode” and deciding whether to create ui_preference, display_mode, prefers_dark_mode, or dark_mode_enabled), drift occurs at the schema-definition level rather than the extraction level. Knowledge Objects address this by defining the controlled vocabulary in advance as part of system design, making predicate selection a lookup rather than a generation task.
5.8 Experiment 8: KO Retrieval at Scale
The preceding experiments used synthetic corpora with at most a few hundred facts. To validate that Knowledge Objects are practical beyond synthetic benchmarks, we evaluated retrieval on 16,309 facts extracted from Wikipedia articles and compared KO performance against Modern Hopfield Networks, the mathematical foundation underlying transformer attention and fast-weight memory systems (Ramsauer et al., 2021).
Methodology.
We constructed a corpus of 16,309 Knowledge Objects by extracting structured facts from randomly sampled Wikipedia articles using our KO compiler. The corpus spans 82 unique predicate types with semantic density $\rho\approx 0.10$ (mean pairwise cosine similarity of embeddings). We evaluated three corpus representations: structured (subject-predicate-object facts), raw text (original Wikipedia sentences), and hybrid (both concatenated). For each representation, we generated 500 test queries using natural language templates and measured top-$k$ retrieval accuracy. To ensure statistical robustness, we ran all experiments with 5 random seeds and report mean $\pm$ standard deviation. The Hopfield baseline implements Modern Hopfield Networks with softmax attention and tests multiple inverse temperature values ($\beta\in\{0.5,1.0,2.0,4.0,8.0\}$).
[…]
System Top-1 Top-5 Top-10 MRR
KO (structured corpus) 45.7% $\pm$ 2.2% 77.6% $\pm$ 1.9% 87.8% $\pm$ 1.4% 0.600 $\pm$ 0.019
KO (hybrid corpus) 33.3% $\pm$ 3.0% 64.8% $\pm$ 3.5% , ,
KO (raw text corpus) 4.1% $\pm$ 1.0% 18.5% $\pm$ 1.6% , ,
Hopfield ($\beta=1$) 0.0% $\pm$ 0.0% 0.0% $\pm$ 0.0% 0.0% $\pm$ 0.0% ,
Hopfield ($\beta=8$) 0.1% $\pm$ 0.2% 0.1% $\pm$ 0.2% 0.1% $\pm$ 0.2% ,
Degradation Under Scale.
A critical question for practical deployment is how retrieval accuracy changes as corpus size increases. Figure 7 and Table 11 reveal that both KO embedding retrieval and Hopfield networks degrade with scale, but at fundamentally different rates. At small scales ($N\leq 100$), both systems achieve near-perfect accuracy: KO reaches 98% and Hopfield ($\beta=8$) reaches 97%. As $N$ increases, Hopfield collapses rapidly, falling to 43% at $N=500$, 10.5% at N=1, [^0] N=1{,}000, and effectively 0% by N=5, [^0] N=5{,}000. KO embedding retrieval degrades more gradually, maintaining 90.5% at $N=500$, 88% at N=1, [^0] N=1{,}000, and reaching 45.7% $\pm$ 2.2% at N=16,309N=16{,}309 in our rigorous 5-seed evaluation. The gap between the two systems widens with scale, reaching approximately 46 percentage points at N=16,309N=16{,}309.
Table 11: Degradation comparison: Both systems degrade with scale, but KO degrades gradually while Hopfield collapses catastrophically. Final row shows rigorous 5-seed result; other rows from single-seed comparison for curve shape.
N KO Embedding Hopfield ($\beta=8$) Hopfield ($\beta=1$)
10 100% 100% 10%
100 98% 97% 1%
500 90.5% 43% 0%
1,000 88% 10.5% 0%
5,000 67% 0% 0%
16,309 45.7% $\pm$ 2.2% 0.5% 0%
5-seed rigorous evaluation; other rows single-seed for degradation curve shape.
Effect of Corpus Representation: The Central Finding.
[…]
Implications for Production Systems.
Production memory systems like ChatGPT Memory and Claude Memory store facts as unstructured text, for example, “User mentioned they like Python” or “User is allergic to shellfish.” This storage format resembles our “raw text” condition rather than our “structured” condition. Our results suggest that as these systems accumulate thousands of user facts, retrieval accuracy may degrade substantially. At N=16,309N=16{,}309 facts, raw text retrieval achieves only 4.1% accuracy compared to 45.7% for structured facts, a gap that would manifest as users experiencing their AI “forgetting” facts that are actually stored but unretrievable. We could not test production memory APIs directly (they are not publicly accessible for programmatic evaluation at scale), but the storage format similarity suggests production systems face comparable degradation. This motivates structured storage as a design principle: the KO compiler’s extraction of subject-predicate-object facts serves a dual purpose of enabling both hash-based lookup and embedding-aligned retrieval.
Comparison with Hopfield.
The contrasting degradation patterns of KO and Hopfield networks provide empirical support for the Orthogonality Constraint. Ramsauer et al. (Ramsauer et al., 2021) proved that transformer attention is mathematically equivalent to a Modern Hopfield Network update. Our results show that Hopfield networks can achieve high accuracy at small scales, 97% at $N=100$ with $\beta=8$, but collapse rapidly as semantic interference accumulates, falling to 10.5% at N=1, [^0] N=1{,}000 and effectively 0% by N=5, [^0] N=5{,}000. KO embedding retrieval, which stores facts at discrete addresses and retrieves via the same cosine similarity mechanism, degrades at roughly half the rate: from 98% at $N=100$ to 45.7% $\pm$ 2.2% at N=16,309N=16{,}309. The key difference is that KO stores facts in separate memory slots (preventing write-time interference), while Hopfield superimposes all patterns in shared continuous parameters where they interfere during both storage and retrieval. This may explain why Hopfield’s collapse is catastrophic rather than gradual: once interference exceeds a threshold, the softmax attention spreads across many similar patterns and retrieval becomes effectively random.
Cost Projections at Enterprise Scale.
[…]
Table 12: Cost and latency projections at enterprise scale (validated to NN=500K for hash, NN=100K for embedding). KO cost includes embedding API (\$<<0.0001/q), vector search (amortized infra), and LLM call with top-10 facts.
N Hash Latency Embed Latency KO Cost Context Cost Ratio
[…]
Projected with ANN index (HNSW). Brute-force would be 300ms and 3s respectively.
KO cost breakdown: embedding query (\$0.0001), vector DB lookup (\$0.0001 amortized), LLM with 10 retrieved facts (\$0.0013). Vector DB assumes \$70/month hosted service at 1M+ queries/month; self-hosted reduces this further.
Figure 7: Degradation Under Scale: KO vs Hopfield. Both KO embedding retrieval and Hopfield ($\beta=8$) achieve near-perfect accuracy at small scales ($N\leq 100$). As corpus size increases, Hopfield collapses catastrophically, falling from 97% at $N=100$ to effectively 0% by N=5, [^0] N=5{,}000. KO degrades more gradually, maintaining approximately 46% at N=16,309N=16{,}309. The roughly 46 percentage point gap at scale suggests that discrete storage degrades gracefully while neural memory fails catastrophically under semantic density. Hopfield with $\beta=1$ (dotted line) fails immediately, achieving only 10% even at $N=10$.
5.9 Experiment 8: Real-World Validation
The preceding experiments used synthetic fact corpora designed to exhibit controlled semantic density. A natural question is whether real-world data exhibits comparable density and produces similar collapse patterns. We validate the Orthogonality Constraint on two real-world datasets spanning different modalities: (1) the UCI Wine Quality dataset containing physicochemical measurements, and (2) CIFAR-100 images encoded with CLIP. Together, these experiments demonstrate that the constraint is not an artifact of synthetic text but a fundamental property of embedding geometry across modalities.
5.9.1 Scientific Measurements (UCI Wine Quality)
Scientific measurements represent a critical use case for AI memory systems: laboratory assistants must store and retrieve precise experimental values (“The pH of sample_001 is 3.21”) without confusing them with similar measurements (“The pH of sample_002 is 3.34”). This is exactly the scenario where neural memory should fail catastrophically: the sentences are nearly identical, differing only in sample identifiers and numeric values.
Methodology.
The UCI Wine Quality dataset (Cortez et al., 2009) contains 11 physicochemical measurements per wine sample: fixed acidity, volatile acidity, citric acid, residual sugar, chlorides, free sulfur dioxide, total sulfur dioxide, density, pH, sulphates, and alcohol content. We converted each measurement to a natural language fact following the template “The [measurement] of [sample_id] is [value] [unit]” (e.g., “The pH of wine_00001 is 3.210”), yielding 64,970 total facts from 6,497 wine samples. We computed same-property semantic density $\rho_{\text{same}}$ as the average pairwise cosine similarity among facts sharing the same measurement type (e.g., all pH measurements), this represents the worst-case interference scenario. We ran 3 trials per $N$ value using all-MiniLM-L6-v2 embeddings.
Table 13: Scientific Measurements (UCI Wine Quality): Neural collapse under extreme semantic density ($\rho_{\text{same}}\approx 0.96$). At N=10, [^0] N=10{,}000, neural memory achieves 0.02% accuracy, effectively zero.
N KO (Hash) KO (Embed) Neural (LAM) $\rho_{\text{same}}$
10 100% 100% 13.3% 0.964
100 100% 87.7% 2.3% 0.960
1,000 100% 57.5% 0.17% 0.960
10,000 100% 28.0% 0.02% 0.961
Results.
The results reveal the most extreme semantic density we have measured: $\rho_{\text{same}}\approx 0.96$, meaning that facts sharing a measurement type (e.g., all pH values across different wines) have 96% identical embeddings. This is substantially higher than our synthetic corpora ($\rho\approx 0.70$) or typical text ($\rho\approx 0.30$$0.50$). Neural memory achieves only 13.3% accuracy at $N=10$ facts, already near random chance, and collapses to 0.02% (effectively zero) at N=10, [^0] N=10{,}000. Even embedding-based semantic search degrades from 100% to 28% under this density. Hash-based retrieval maintains 100% accuracy throughout.
5.9.2 Visual Data (CIFAR-100 + CLIP)
To confirm the Orthogonality Constraint applies beyond text, we evaluated memory systems on visual data using CIFAR-100 (Krizhevsky, 2009), which contains 10,000 test images across 100 fine-grained categories. Images were encoded using CLIP ViT-B/32 (Radford et al., 2021) ($d=512$).
Methodology.
We measured image-to-image (I2I) retrieval: given an image embedding as query, retrieve that exact image from the store. We computed overall semantic density $\rho$ as the average pairwise cosine similarity across all stored images, and intra-class density $\rho_{\text{intra}}$ as the average similarity among images of the same category (e.g., all “apple” images). We ran 3 trials per $N$ value.
Table 14: Image Retrieval (CIFAR-100 + CLIP): Neural collapse under visual semantic density. Intra-class $\rho\approx 0.82$ indicates images of the same category are nearly indistinguishable to CLIP.
N KO (Hash) KO (I2I) Neural (I2I) $\rho$ $\rho_{\text{intra}}$
10 100% 100% 20.0% 0.74 0.27
50 100% 100% 3.3% 0.74 0.83
250 100% 100% 0.4% 0.74 0.82
1,000 100% 100% 0.1% 0.74 0.83
2,000 100% 100% 0.05% 0.74 0.82
Results.
Visual semantic density reaches $\rho=0.74$ overall and $\rho_{\text{intra}}=0.82$ within categories, images of the same class are 82% similar to CLIP. Neural memory collapses from 20% at $N=10$ to 0.05% at N=2, [^0] N=2{,}000. Hash-based retrieval maintains 100% throughout.
5.9.3 Unified Analysis
Table 15 synthesizes results across modalities and compares semantic density levels.
Table 15: Real-world validation summary: The Orthogonality Constraint applies across modalities. Higher $\rho$ produces faster collapse. Only hash-based retrieval survives.
Dataset Modality $\rho$ Neural @ N=1K KO Hash
Synthetic facts Text 0.70 1.3% 100%
CIFAR-100 + CLIP Images 0.82 (intra) 0.1% 100%
UCI Wine Quality Scientific text 0.96 0.17% 100%
These results establish three key findings:
  1. 1.
    Real-world data exhibits extreme semantic density. Scientific measurements reach $\rho=0.96$, higher than any synthetic corpus we could construct. Visual data exhibits $\rho_{\text{intra}}=0.82$ within categories.
  2. 2.
    The Orthogonality Constraint is modality-agnostic. Whether the embedding space encodes text (sentence transformers) or images (CLIP), the same geometric interference produces the same collapse pattern.
  3. 3.
    Only discrete addressing survives. Hash-based retrieval maintains 100% accuracy across all datasets and scales. Embedding-based retrieval degrades under density; neural parameter storage collapses catastrophically.
Implications.
These findings have direct implications for AI systems in high-precision domains. A laboratory assistant that cannot distinguish measurements from different samples is worse than useless. A medical AI that confuses patient vitals across visits poses safety risks. A financial system that retrieves the wrong transaction amounts creates liability. The Orthogonality Constraint is not an abstract theoretical concern but a practical barrier to deploying neural memory wherever structurally similar facts must be stored and retrieved without confusion. Knowledge Objects, with hash-based addressing derived from structured identifiers, provide the interference-free storage these applications demand.
@@ 6 Discussion
Neural memory systems that store facts in shared continuous parameters will experience interference proportional to semantic density, and only systems employing discrete addresses can guarantee zero interference between stored facts. This finding does not represent a limitation of current models that might be resolved through scaling or architectural refinements, rather, it reflects a geometric constraint inherent to continuous vector storage that no amount of engineering can circumvent. When two embedding vectors occupy nearby regions of the space, any linear combination that retrieves one will partially retrieve the other; this is not a bug but a mathematical property of inner products.
@@
Current AI systems attempt to accomplish everything within the “neocortex”, that is, neural weights and attention mechanisms, which explains why they fail at precise factual storage despite succeeding at pattern recognition, generation, and reasoning. Our Knowledge Objects implement the missing “hippocampus” in this analogy: not as a biological simulation, but as the engineering solution to the same computational problem that evolution solved through anatomical specialization.
A Neurosymbolic Architecture.
The hybrid architecture we propose, discrete symbolic memory interfacing with neural generation, instantiates a neurosymbolic design in which each subsystem handles the computations it performs best. Knowledge Objects provide the symbolic component: explicit structure (subject-predicate-object facts), discrete addressing (hash-based lookup), deterministic operations (surgical updates, version chains), and schema constraints (controlled vocabularies). The LLM provides the neural component: pattern completion, semantic similarity, natural language understanding, and flexible generation. The learned router mediates between subsystems, classifying queries by structure rather than content.
This division mirrors the System 1/System 2 framework from cognitive science (Kahneman, 2011): the neural pathway offers fast, intuitive responses for open-ended queries, while the symbolic pathway provides slow, deliberate retrieval for precise factual questions. Unlike purely neural approaches that attempt both functions in shared continuous parameters, or purely symbolic approaches that lack the flexibility of learned representations, the neurosymbolic design allocates each function to the appropriate substrate. The Orthogonality Constraint provides theoretical justification for this separation: symbolic storage is not merely convenient but necessary for interference-free episodic memory at scale.
@@
Evaluating Production Systems.
The Orthogonality Constraint provides a framework for evaluating production memory systems. Both ChatGPT Memory and Claude Memory have adopted discrete external storage, ChatGPT stores explicit facts as short timestamped entries (e.g., “User is allergic to shellfish”), while Claude stores information in Markdown files with on-demand retrieval tools. This architectural choice is correct: discrete storage escapes the geometric interference that makes neural episodic memory impossible at scale. However, both systems remain incomplete. They store unstructured text extracted by LLMs:
  • Current: Unstructured text (“User mentioned they like Python”)
  • KO: Structured facts with controlled vocabulary: (user, preferred_language, Python)
This distinction matters. Unstructured storage lacks key-based identity (preventing deterministic lookup), schema constraints (causing predicate drift), and version semantics (causing ambiguity about current values). Our multi-model comparison demonstrates that schema drift and version ambiguity affect all tested frontier models, confirming these are architectural limitations of the unstructured approach rather than bugs in any particular implementation. Knowledge Objects provide the structured semantics that production systems lack: hash-based addressing for deterministic lookup and surgical updates, controlled vocabularies for schema consistency, and version chains for explicit supersession. The path from current production systems to reliable episodic memory requires not just discrete storage (which they have) but structured discrete storage (which they lack).
@@
Our analysis has several limitations that should inform interpretation of the results. First, our neural baseline (LAM) is a simplified proxy for production systems that store facts via superposition; we treat it as a mechanistic lower bound that isolates the interference mechanism arising whenever semantically overlapping keys are stored in shared continuous parameters. Production implementations may include mitigations such as gating and regularization that shift the collapse curve rightward, but absent explicit pattern separation, overlap-driven interference persists when keys are semantically clustered, our claim concerns the existence of interference under semantic density, not the performance of any specific implementation. Second, our experiments use a single embedding model (all-MiniLM-L6-v2 with $d=384$ dimensions); different embedding dimensions would shift the quantitative results, though the qualitative finding, that collapse occurs at $N$ far below theoretical capacity, should hold. Third, we measure retrieval accuracy on synthetic fact corpora, not end-to-end task performance on downstream applications, though retrieval accuracy is a prerequisite for task success. Fourth, our multi-model comparison tests only three models from two providers; additional models might reveal different failure patterns, though the consistency we observe suggests these are general phenomena. Fifth, our experiments focus exclusively on textual facts; we discuss generalization to other modalities below but do not empirically validate multimodal performance.
Generalization to Multimodal Systems.
While our main experiments focus on textual facts, the Orthogonality Constraint applies to any embedding space. We validate this in Section 5 using CIFAR-100 images encoded with CLIP ViT-B/32 (Radford et al., 2021). Visual semantic density reaches $\rho=0.74$ overall and $\rho=0.82$ intra-class (images of the same category), causing neural memory to collapse from 20% at $N=10$ to 0.05% at N=2, [^0] N=2{,}000. Hash-based retrieval maintains 100% accuracy throughout, confirming that the constraint is modality-agnostic.
In domains such as medical imaging or satellite imagery, visual semantic density may be particularly high: images that appear similar (two chest X-rays, two aerial photographs of farmland) but represent distinct entities will occupy nearby regions of embedding space, producing the same retrieval interference we document for text and validate for natural images. Knowledge Objects generalize naturally to multimodal content: the object field holds arbitrary media addressed by content hash, with modality-specific embeddings enabling retrieval while hash-based addressing guarantees zero storage interference. The neurosymbolic architecture thus extends beyond language to any domain where precise retrieval of specific instances matters.